Solve the system: 2x + 3y = 6, x - 3y = 8
Thank you!

Answer by stanbon(75887)

Solve the system:
2x + 3y = 6
x - 3y = 8
---------------
Add the two equations to get:
3x = 14
x = 14/3
----
Sustitute for "x" and solve for "y":
(14/3) - 3y = 8
3y = (14/3)-(24/3)
3y = -10/3
y = -10/9
================
solution: (14/3 , -10/9)
============================
Cheers,
Stan H.